A glider with mass sits on a frictionless horizontal air track, connected to a
spring with force constant k = 5.00 N/m. You pull on the glider, stretching
the spring 0.100 m, and release it from rest. The glider moves back toward
its equilibrium position (x=0). What is its x-velocity when x = 0.080 m?
1
Expert's answer
2020-04-15T10:40:16-0400

From the conservation of energy:


"0.5mv^2=0.5kA^2-0.5kx^2"

"(0.2)v^2=(5)(0.1^2-0.08^2)"

"v=0.3\frac{m}{s}"